OpenStudy (anonymous):

If f(x) = + 8, what is f(x) when x = 10?

Mehek (mehek14):

you mean f(x) = x + 8?

OpenStudy (anonymous):

If it was f(x) = x - 2 and x was 10: f(10) = (10) - 2 f(10) = 8 If it was f(x) = - 2 and x was 10: f(10) = -2

OpenStudy (anonymous):

actually it was supposed to me x over 2

Mehek (mehek14):



OpenStudy (anonymous):

yes

Mehek (mehek14):

ok so